OpenStudy (anonymous):

Let f(x)=3x-6 and g(x)=x-2 find f/g and it's domain

OpenStudy (anonymous):

f/g = (3x-6) / (x-2) the domain is all x but the x's that make the denominator zero so the domain is : all x except 2

OpenStudy (anonymous):

f/g = 3

OpenStudy (anonymous):

(it is simplified like this) but still x cant be 2.

OpenStudy (anonymous):

there is a "hole" at x=2

OpenStudy (anonymous):

and f/g can not be 3

OpenStudy (anonymous):

yes it can be 3.. it is 3 everywhere except at x=2.

OpenStudy (anonymous):

\[f/g=(3x-6)\div(x-2)=3\] then we cannot find x for which f/g=3

OpenStudy (anonymous):

when someone tells you y=2 you know that y is 2 for every x. how would you graph that ?

OpenStudy (anonymous):